# AR-TTA: A Simple Method for Real-World Continual Test-Time Adaptation

### v1

@@ 3.3 Dynamic Batch Norm Statistics
Batch normalization [16] (BN) was created for reducing the internal covariate shift occurring during model training. It normalizes the distribution of a batch of input data utilizing the calculated running statistics, namely mean and variance, which is saved after the model training and used for the inference at test-time. While testing on out-of-distribution data, saved statistics are not correct and the normalization process fails to produce data with standard normal distribution which leads to poor model performance. Therefore, state-of-the-art test-time adaptation methods [39, 29, 37, 30] usually discard statistics calculated during training and estimate data distribution based on each batch of data separately. However, this way of estimating the statistics is flawed, since the sample size from data is usually too small to correctly estimate the data distribution, depending on batch size. Furthermore, samples might be temporally correlated (e.g. video input), which also is the cause of incorrect statistics estimates. In such cases, BN statistics from source data might be useful and closer to the actual data distribution, compared to the estimated values. To robustly estimate the correct normalization statistics we take the inspiration from [15] and propose to estimate BN statistics $\phi_{t}=(\mu_{t},\sigma_{t})$ at time step $t$ during test-time by linearly interpolating between saved statistics from source data $\phi^{S}$ and calculated values from current batch $\phi^{T}_{t}$:
@@
where $\beta_{ema}$ is a parameter that weights the influence of saved and currently calculated statistics.
Since the severity of distribution shift might vary, we need to adequately adjust the value of $\beta_{ema}$. It should be large in cases when the distribution shift is severe compared to the source domain and low when the distributions are similar. Following [15], we utilize the symmetric KL divergence as a measure of distance between distributions $D(\phi_{t-1},\phi_{t}^{T})$:
@@
where $\gamma$ is a scale hyperparameter.
To compensate for the fact that the current distribution can be wrongly estimated and to provide more stability for the adaptation, we take into account previous $\beta_{1:t-1}$ values and use an exponential moving average for $\beta_{ema}$ update:
@@
where $\alpha$ is a hyperparameter.
The difference between our method and MECTA [15] is that we do not use the exponential moving average of calculated BN statistics, but instead, we keep the statistics from source data intact. We are motivated by the fact that changing this value cause the inevitable forgetting, accumulation of BN statistics estimate error, and susceptibility to temporal correlation. By keeping this value constant, we make sure that the estimation of the statistics on every batch does not cause the degradation of performance on domains similar to the source domain, leading to worse results than the frozen source model and abandonment of the legitimacy of using TTA methods. At the same time, we allow for a slight drift away from source data statistics, by using the exponential moving average of $\beta$ parameter, giving enough flexibility for the adaptation for severe domain shifts.
@@ 4 Experiments
Datasets and Benchmarks
We evaluate the methods on four image classification tasks: CIFAR10-to-CIFAR10C, ImageNet-to-ImageNet-C, CLAD-C continual learning benchmark [36] adapted to the test-time adaptation setting, and the benchmark created from the SHIFT dataset [34].
CIFAR10-to-CIFAR10C and ImageNet-to-ImageNet-C are widely used tasks in TTA. They involve training the source model on train split of clean CIFAR10/ImageNet datasets [19, 8] and test-time adaptation on CIFAR10C/ImageNet-C. CIFAR10C and ImageNet-C consist of images from clean datasets which were modified by 15 types of corruptions with 5 levels of severity[14]. They were first used for evaluating the robustness of neural network models and are now widely utilized for testing the adaptation capabilities of TTA methods. We test the adaptation on a standard sequence of the highest corruption severity level 5, frequently utilized by previous approaches [39, 29, 9]. For ImageNet-C we utilize a subset of 5000 samples for each corruption, based on RobustBench library [7], following [39].
CLAD-C [36] is an online classification benchmark for autonomous driving with the goal to introduce a more realistic testing bed for continual learning. Even though, to the best of our knowledge, it has not yet been used for testing TTA before, we chose this benchmark inspired by its realistic nature and the goal to simulate more real-world application setup. It consists of natural, temporal correlated, and continuous distribution shifts created by utilizing the data from SODA10M dataset [13]. The images taken at different locations, times of day, and weathers, are chronologically ordered, inducing distribution shifts in labels and domains. The classification task was created by cutting out the annotated 2D bounding boxes of six classes and using them as separate images for classification. Since it is designed for testing the continual learning setup and the model is originally supposed to be trained sequentially on the train sequences, we slightly modify the setup and pre-train the source model on the first train sequence. TTA is continually tested on the 5 remaining ones with the total number of 17092 images.
The SHIFT dataset is a synthetic autonomous driving dataset designed for continuous multi-task domain adaptation. It provides multiple types of data from the perspective of a car, including RGB images with various types of annotations. Images are taken in numerous types of realistic domains simulated in a virtual environment, including different weather and times of day. This dataset is not designed for a classification task. However, satisfied with the amount of data and the realism of the dataset, we decided to create the classification setup, using the same procedure as for the CLAD-C dataset, utilizing the 2D bounding box annotations. We train the source model on images taken at clean weather during the day and test the adaptation methods for various weather combinations and times of day. We end up with 14 different domains resulting in total of 380667 images. The high number of test samples provides a good simulation of what could happen during model deployment in a real world, ever-changing environment. We called our original continual test-time adaptation benchmark on SHIFT dataset SHIFT-C.
Methodology
@@
Taking into account the realistic scenario, we examine the continual test-time adaptation setup where the model is continually adapted to new domains without any weight resetting to the source model state, unless it is a part of a tested method. To simulate the continuous stream of data and the need for the model to adapt quickly to the data it is provided with, we use a low batch size of 10. This is also the batch size commonly used in the online continual learning [25]. Considering a practical application on embedded device, low batch size could be also a result of limited computation resources and memory constraints, making it impossible to use higher batch sizes. We believe that it is important and more practical for TTA methods to be able to work on a reduced number of samples.
We primarily assess the methods using a straightforward mean classification accuracy metric. Additionally, to make every class equally important in class-imbalanced datasets, we use the average mean class accuracy (AMCA) metric. This metric calculates the mean accuracy over all classes, averaged for each domain.
The results are averaged between 3 random seeds. Samples from CIFAR10C and ImageNet-C are shuffled. Considering the sequential nature of data in CLAD-C and SHIFT-C benchmarks (video sequences), we did not want to shuffle images. Instead, to get seed-averaged results, we trained 3 source models with 3 different seeds and averaged the results between experiments with different models.
Baselines
To evaluate the performance of our method and validate its efficacy in handling realistic domain shifts, we conduct experiments involving five state-of-the-art methods as baselines: TENT-continual [37], EATA [29], CoTTA [39], and SAR [30]. Moreover, we show results for discarding BN statistics from source data and calculating the statistics for each batch separately (BN stats adapt) [33]. Additionally, we showcase the results obtained from the frozen source model to verify the effectiveness of adaptation (Source).
To provide a fair comparison with methods that do not require saving a small source data memory bank, we also show the results of our proposed method without the usage of replay memory. Furthermore, we show the results of baselines with a simple replay strategy method added in the appendix.
Implementation Details
Following other state-of-the-art TTA methods, we use pre-trained WideResnet28/ResNet50 models from RobustBench [7] model zoo for CIFAR10-to-CIFAR10C/ImageNet-to-ImageNet-C tasks. On the rest of the benchmarks, we utilize ResNet50 architecture with weights pre-trained on ImageNet obtained from torchvision library [26] and finetuned to the source data for the specific benchmark. Images from CLAD-C and SHIFT-C benchmarks are resized to 224x224 before being processed by the network. For our method, we use an SGD optimizer with a momentum equal to 0.9. The learning rate is set to 0.00025 for ImageNet-C and 0.001 for the rest of the benchmarks. The default replay memory size is 2000 samples, which is commonly used in continual learning settings [27] and adds only minor storage requirements. We provide experiments with different memory sizes in the appendix. The $\gamma$ value in Equation 7 is set to 10 for CIFAR10C and ImageNet-C, and 0.1 for CLAD-C and SHIFT-C. The $\alpha$ value for the exponential moving average of $\beta_{ema}$ in Equation 8 is set to 0.2. The initial $\beta_{ema}$ is equal to 0.1. The parameters $\psi$ and $\rho$ of the beta distribution, which are utilized to sample the interpolation parameter $\lambda$ for mixup augmentation (Equations 34), are both set to a value of 0.4. We provide the results for different shapes of beta distribution in the appendix.
Since the batch size in our experiments is significantly lower than the ones used in the previous TTA works and we use two datasets on which the methods were not tested before, to allow for a fair comparison, we choose the learning rate for each method based on a grid search. The optimizers are chosen following the methods’ papers. For datasets that were not used in an EATA [29] paper, we also searched for the most optimal value of EATA’s method parameter $\epsilon$, which is responsible for filtering the redundant samples for model adaptation. The description and all the results of the parameter search can be found in the appendix. For the ease of prototyping and testing, we utilized the Avalanche library for continual learning [23].
@@ 4.1 Results
Artificial Domain Shifts

### v2

@@ 3.3 Dynamic Batch Norm Statistics
Batch normalization [Ioffe and Szegedy(2015)] (BN) layers normalize network values using running statistics from initial training data, applied during test-time. However, out-of-distribution data can disrupt this process, leading to poor model performance. Test-time adaptation methods [Wang et al.(2022)Wang, Fink, Gool, and Dai, Niu et al.(2022)Niu, Wu, Zhang, Chen, Zheng, Zhao, and Tan, Wang et al.(2021)Wang, Shelhamer, Liu, Olshausen, and Darrell, Niu et al.(2023)Niu, Wu, Zhang, Wen, Chen, Zhao, and Tan] often recalculate statistics per each batch separately. However, this approach can be flawed due to small sample sizes and temporal data correlation. In these cases, using BN statistics from source data may offer a more accurate distribution estimation.
To perform a robust statistics estimation, we take the inspiration from [Hong et al.(2023)Hong, Lyu, Zhou, and Spranger] and propose to calculate BN statistics $\phi_{t}=(\mu_{t},\sigma_{t})$ at time step $t$ during test-time by linearly interpolating between saved statistics from source data $\phi^{S}$ and calculated values from current batch $\phi^{T}_{t}$:
@@
where $\beta_{ema}$ is a parameter that weights the influence of saved and currently calculated statistics. Since the severity of the distribution shift might vary, we need to adequately adjust the value of $\beta_{ema}$. It should be large in cases when the distribution shift is severe compared to the source domain and low when the distributions are similar. Following [Hong et al.(2023)Hong, Lyu, Zhou, and Spranger], we utilize the symmetric KL divergence as a measure of distance between distributions $D(\phi_{t-1},\phi_{t}^{T})$:
@@
where $\gamma$ is a scale hyperparameter. To compensate for the fact that the current distribution can be wrongly estimated and to provide more stability for the adaptation, we take into account previous $\beta_{1:t-1}$ values and use an exponential moving average for $\beta_{ema}$ using $\alpha$ coefficient:
@@
Our method differs from MECTA [Hong et al.(2023)Hong, Lyu, Zhou, and Spranger] by preserving the original BN statistics instead of using exponential moving averages. We are motivated by the fact that changing this value causes the inevitable forgetting, and susceptibility to temporal correlation. We ensure consistent performance on similar domains to the source data. At the same time, we allow for a slight drift away from source data statistics, by using the exponential moving average of $\beta$ parameter, giving enough flexibility for the adaptation to severe domain shifts.
@@ 4 Experiments
Datasets and Benchmarks. We evaluate the methods on multiple image classification datasets with domain shifts. We use CIFAR10C and ImageNet-C as benchmarks with artificial, corruption-based shifts, widely utilized for evaluation of the test-time adaptation approaches [Niu et al.(2023)Niu, Wu, Zhang, Wen, Chen, Zhao, and Tan, Wang et al.(2022)Wang, Fink, Gool, and Dai, Niu et al.(2022)Niu, Wu, Zhang, Chen, Zheng, Zhao, and Tan, Döbler et al.(2022)Döbler, Marsden, and Yang, Wang et al.(2021)Wang, Shelhamer, Liu, Olshausen, and Darrell]. The experiments with natural shifts include tests on CIFAR10.1 [Recht et al.(2018a)Recht, Roelofs, Schmidt, and Shankar] and SHIFT [Sun et al.(2022)Sun, Segu, Postels, Wang, Van Gool, Schiele, Tombari, and Yu] datasets, and CLAD-C continual learning benchmark [Verwimp et al.(2023)Verwimp, Yang, Parisot, Hong, McDonagh, Pérez-Pellitero, De Lange, and Tuytelaars] adapted to the test-time adaptation setting.
CIFAR10.1 was designed to minimize the distribution shift relative to the original CIFAR10 dataset. This allows us to test the robustness of TTA methods to very delicate distribution shifts. The goal of CLAD-C benchmark was to introduce a more realistic testing bed for continual learning. The images taken at different locations and conditions are chronologically ordered, inducing distribution shifts in labels and domains. It consists of 6 test sequences. We pre-train the source model on the first train sequence. TTA is continually tested on the 5 remaining ones with a total number of 17092 images. In the SHIFT dataset, images are taken in numerous types of realistic domains simulated in a virtual environment, including different weather and times of day. We consider it a natural domain shift benchmark due to its realistic representation of distribution shifts, which are induced by real conditions rather than image corruptions. We train the source model on images taken in clear weather during the day and test the adaptation methods for various weather combinations and times of day. We end up with 14 different domains resulting in a total of 380667 images. We called our original continual test-time adaptation benchmark the SHIFT-C. More details can be found in the appendix.
@@
Methodology. We examine the continual test-time adaptation setup where the model is continually adapted to new domains without resetting the model unless it is part of a tested method. To simulate the continuous stream of data and the need for the model to adapt quickly, we use a low batch size of 10. This is also the batch size commonly used in the online continual learning [Mai et al.(2022)Mai, Li, Jeong, Quispe, Kim, and Sanner]. In embedded applications, limited resources often mean smaller batch sizes, necessitating TTA methods to work effectively with fewer samples. We assess the methods using mean classification accuracy and average mean class accuracy (AMCA). The latter calculates the mean accuracy over all classes averaged for each domain, making every class equally important in class-imbalanced datasets.
Baselines. We conduct experiments involving six state-of-the-art methods as baselines: TENT-continual [Wang et al.(2021)Wang, Shelhamer, Liu, Olshausen, and Darrell], EATA [Niu et al.(2022)Niu, Wu, Zhang, Chen, Zheng, Zhao, and Tan], CoTTA [Wang et al.(2022)Wang, Fink, Gool, and Dai], SAR [Niu et al.(2023)Niu, Wu, Zhang, Wen, Chen, Zhao, and Tan], and RMT [Döbler et al.(2022)Döbler, Marsden, and Yang]. Moreover, we show results for discarding BN statistics from source data and calculating the statistics for each batch separately (BN-1) [Schneider et al.(2020)Schneider, Rusak, Eck, Bringmann, Brendel, and Bethge]. We also report the performance of the frozen source model (Source). To provide a fair comparison, the results without a replay strategy of RMT and our method (AR-TTA) are presented.
Implementation Details. Following other state-of-the-art TTA works, we use pre-trained WideResnet28/ResNet50 models from RobustBench [Croce et al.(2020)Croce, Andriushchenko, Sehwag, Debenedetti, Flammarion, Chiang, Mittal, and Hein] model zoo for CIFAR10C and CIFAR10.1/ImageNet-C datasets. On the rest of the benchmarks, we utilize ResNet50 architecture with weights pre-trained on ImageNet obtained from torchvision library [maintainers and contributors(2016)] and finetuned to the source data for the specific benchmark.
Our method utilizes the default replay memory size of 2000 samples, which is commonly used in continual learning settings [Masana et al.(2022)Masana, Liu, Twardowski, Menta, Bagdanov, and van de Weijer] and adds only minor storage requirements. The RMT method utilizes all of the source training data, following the original work. The rest of the details regarding the technical side of experiments can be found in the appendix.
@@ 4.1 Results
The results for all of the tested benchmarks are presented in Table 1. Our method can achieve the best average accuracy of 75.5% indicating the robustness of our approach. Moreover, it shows a solid performance even without the replay method.